OpenStudy (anonymous):

what value, in place of the question mark, makes the polynomial below a perfect square trinomial? ... x^2+14x+ ?

OpenStudy (mathstudent55):

Look at a^2 + 2ab + b^2 = (a + b)(a + b) Now copmpare with what you have: a^2 = 1 so a = 1 The middle term is 2ab, and you have 14x, so 2abx = 14x, but a = 1 so 2b = 14 b = 7 The last term of the trinomial is b^2, and b = 7, so b^2 = ______

OpenStudy (anonymous):

is it 49, lol

OpenStudy (mathstudent55):

right

OpenStudy (anonymous):

okay, thanks (: